Grade 11Mechanics

a particle starting from rest its acceleration at an instant t is a=10/V+1, where V is instantaneous velocity . the distance travelled after attaining a velocity of 15 m/s is

Akshay keshwani
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear Akshay
 
a = v dv/ds
 
Now
10 / (v+1) = v dv/ds
 
Now
 
10 ds = v (v+1) dv
 
on integrating
 
10 s = v^3/3  + v^2/ 2  + c
 
at v=0
 
s=0
 
hence
10 s = v^3/3  + v^2/ 2  
 
 
now if v = 15 m/sec
 
10s = 3375/3 + 225/2
 
10s = 1125 + 112.5
 
10s = 1237.5
 
s = 123.75 m